### v3

# A Cognitive Evaluation Benchmark of Image Reasoning and Description for Large Vision-Language Models

Abstract
Large Vision-Language Models (LVLMs), despite their recent success, are hardly comprehensively tested for their cognitive abilities. Inspired by the prevalent use of the “Cookie Theft” task in human cognition test, we propose a novel evaluation benchmark to evaluate high-level cognitive ability of LVLMs using images with rich semantics 111 The code and the data of CogBench is released at: https://github.com/xiujiesong/Cog-Bench. . It defines eight reasoning capabilities and consists of an image description task and a visual question answering task. Our evaluation on well-known LVLMs shows that there is still a large gap in cognitive ability between LVLMs and humans.
1 Introduction
Recently, with the emergence of Large Language Models (LLMs) such as GPT-4 [1], the cognitive abilities of language models have reached a new level [2]. They demonstrated remarkable performance in many tasks [3]. In Vision Language (VL), several researchers [4, 5, 6] endeavor to boost Vision-Language Pre-trained Models (VLPMs) by integrating powerful LLMs [7, 8], referred to as Large Vision-Language Models (LVLMs). With LLM as the “brain”, the cognitive abilities of LVLMs are also improved and more challenging tasks, such as table, chart, and document reasoning, can be solved [9]. Some state-of-the-art LVLMs, such as GPT-4V and GPT-4o[1], are progressing towards human-level cognitive abilities [9]. Thus, there’s a growing interest in evaluating the cognitive abilities of LVLMs. Such evaluation usually amounts to co-reasoning of images and text, e.g., question answering about a picture. Though some LVLM evaluation benchmarks, such as MME [10], MMBench [11], and SEED Bench [12], evaluate cognitive reasoning ability as one aspect of their evaluation, they do not provide a comprehensive evaluation of higher-level reasoning abilities and most of the images they use require relatively little reasoning to understand.
In our study, we draw inspiration from the “Cookie Theft” picture description task (Figure 3), a key component of the Boston Diagnostic Aphasia Examination [13], widely utilized in clinical practice within speech-language pathology [14] and cognitive function screening [15]. Notably, despite being designed over half a century ago, this picture remains prevalent in contemporary psychological discussions.
Figure 1: Cookie Theft picture description task. The descriptions in the green frame and the orange frame were respectively produced by a 75-year-old healthy man and a 66-year-old woman with probable Alzheimer’s dementia [^3] The two samples are extracted from DementiaBank (https://dementia.talkbank.org/), which records description transcripts from both healthy subjects and dementia patients..
Figure 2: The comparison between our images and those from the previous visual reasoning tasks. Compared to our images, image (a) has fewer entities and CoRs, image (b) and (c) have some entities, but fewer CoRs.
Is it possible to transfer the success of Cookie Theft picture description in human cognition test to evaluating cognitive abilities of LVLMs? Linguists and psychologists [14] conducted an analysis to determine the factors contributing to the success of “Cookie Theft”. The study revealed that the narrative includes information with varying levels of importance and encompasses a broad range of semantic categories. It was observed that during the description of “Cookie Theft”, individuals with intact cognitive functions exhibit their cognitive prowess by logically deducing the events and their interconnections. In contrast, those with cognitive impairments tend to merely list the superficial aspects of the situation without deeper reasoning. For instance, in Figure 3, by comparing the descriptions produced by the healthy man and the woman with probable Alzheimer’s dementia, we can identify the following differences:
  • The description produced by a healthy man uses “mother” instead of “lady”, indicating the reasoning about character relationship.
  • The healthy man used “stealing cookies” instead of “taking cookies”, indicating his reasoning about this high-level event. The description produced by the patient even did not mention this event at all.
  • The healthy man also used “the mother is preoccupied” and “happy” to describe people’s mental state.
  • The description also reflects the causal relationships between events. Because “the mother is preoccupied” and “not paying attention to her kids”, “kids are stealing cookies.
Through these reasoning processes, the difference of cognitive abilities between the two individuals is reflected in their descriptions. A picture that can evaluate cognitive functions needs to be carefully designed and crafted. Tasnim et al. [16] introduced guidelines for drawing pictures similar to Cookie Theft, which is consistent with findings mentioned above. Generally speaking, compared to ordinary images, Cookie Theft-like images feature i) a prominent story theme, ii) richer content, iii) display complex relationships among entities, and thus require stronger cognitive abilities to understand and describe.
With the above design principles, we propose to construct a Cognitive Evaluation Benchmark, named as CogBench, which consists of high quality Cookie Theft-like images to evaluate the cognitive reasoning ability of LVLMs. CogBench defines eight core cognitive reasoning capabilities, including reasoning about special time, location, character, character relationship, event, event relationship, next moment event and mental state. Both a generative Image Description task and a discriminative Visual Question Answering (VQA) task are designed. Our main contributions are as follows:
  • To the best of our knowledge, this is the first-of-its-kind attempt to incorporate a cognitive test designed for humans into the evaluation of LVLMs.
  • We created and open-sourced the largest VL dataset with Cookie Theft-like images for evaluating LVLM’s cognitive abilities.
  • Our evaluation on existing LVLMs shows that a large gap exists between the cognitive ability of LVLMs and human beings, indicating CogBench will be a valuable evaluation benchmark in the near future.
2 Dataset Construction
In this section, we will introduce the construction of CogBench. We will introduce image collection, annotation, tasks and the statistics of CogBench. Ethical considerations of CogBench are discussed in Appendix G.
2.1 Image Collection
Based on previous studies [14, 16], we set the following image collection criteria:
  • Rule 1: Story-telling The image depicts an interesting story. For instance, the Cookie Theft picture tells the story of a mother busy washing dishes while two kids takes the opportunity to stand on a stool and sneakily steal cookies.
  • Rule 2: Rich Chain-of-Reasonings Images should display rich Chain-of-Reasonings (CoRs) in a scene. A CoR connects low-level observations in an image to produce a high-level reasoning conclusion or connects the cause and effect of events. For example, “The mother is busy washing dishes. + The boy is standing on the stool behind the mother. + The girl standing by the boy is shushing him. + The boy is fetching cookies from the jar in the cabinet. $\rightarrow$ The boy and girl are stealing cookies.” is a CoR about high-level event “stealing cookies”. Note that the story is actually constructed via these CoRs.
  • Rule 3: Content Simplicity Images should contain rich content but not be overly complex. The number of subjects should be limited for better emphasis on the key points of the story.
With the above criteria, we manually collect pictures from Pinterest [^4] https://www.pinterest.com/, and the Cookie Theft is also included in CogBench. Figure 2 shows the differences between our images and those from other datasets by examples.
2.2 Image Annotation
Figure 3: An example of Description task from CogBench.
Human annotators, mostly undergraduate or graduate students aged 18-28, are hired to annotate the collected images [^5] Wages and recruitment process details can be found in Appendix G.. As shown in Figure 3, the annotation includes three parts: [Entities], [CoRs] and [Description]. By annotating [Entities] and [CoRs], we aim to evaluate the low-level recognition ability and high-level cognitive reasoning ability of models respectively based on description. [Description] is annotated as the reference description for the image. The three parts are annotated in that order.
Entity Annotation We ask annotators to list as many entities in the image as possible and entities that are difficult to recognize should be omitted.
CoR Annotation In order to evaluate model cognition in a fine-grained manner, the following eight reasoning capabilities are annotated with CoRs:
  • Special Time Reasoning: reasoning about the special time of the story in the image, e.g., festivals, seasons etc.
  • Location Reasoning: reasoning about the location of the story in the image, e.g., near the school etc.
  • Character Reasoning: reasoning about the character of the subjects in the image, e.g., police officer etc.
  • Character Relationship Reasoning: reasoning about relationship between characters in the image, e.g., “the woman is the mother of the kids.”
  • Event Reasoning: reasoning about high-level events in the current moment and previous moments in the image based on clues in the picture. The difference between high-level and low-level lies in how much semantic information the event contains, e.g., “stealing cookies” is a higher-level event compared to “taking cookies” as it additionally conveys the semantic of “taking advantage without permission or knowledge.”
  • Event Relationship Reasoning: reasoning about causal and temporal relationship between different events in the image. For instance, “the sink is overflowing because the mother left the tap on.”
  • Next Moment Event Reasoning: reasoning about the event that will happen in the next moment. For example, “The police officer will reprimand the boy who violates the rules.”
  • Mental State Reasoning: reasoning about mental state of subjects in the image, e.g., daydreaming etc.
Description Summary Annotators are last asked to write a description to tell the entire story in the image based on the annotated [Entities] and [CoRs].
The complete annotation instruction is in Appendix A. Considering different people may have different understanding about some images, we ask three annotators to annotate each image. Then, we manually merge the three annotations as one final annotation by majority vote. For the [Description], if two or more annotators understand the image in the same way, we accept the best descriptions as the final description. For [Entities] and [CoRs], we first accept most of the entities and CoRs that are annotated by at least two annotators. Other entities and CoRs are also included if they are reasonable. For images with significantly different understandings from the three annotators, we will discard these images.
2.3 Tasks in CogBench
We design a generative Image Description task and a discriminative Multiple-Choice Question Answering task in CogBench.
2.3.1 Image Description Task
This is the primary task of the benchmark. The difference between our description task and existing image description task [17, 18, 19] is that we expect LVLMs understand and describe the story in the image through high-level cognitive reasoning. For instance, in Figure 3, the description of the image should not only include what is in the picture but also focus on elucidating the story of “parents are going out to attend a formal event, and the mother is refusing the boy to accompany them” through a series of reasonings.
2.3.2 Visual Question Answering Task
Figure 4: Generating a MCQ from event reasoning annotations.
The VQA task features standard four-option Multiple-Choice Questions, easing the evaluation process. Like the Description task, VQA questions involve different types of high-level cognitive reasoning, as illustrated by the question about events in Figure 4.
[…]
2.4 CogBench Statistics
Table 1: Distribution of CoRs and questions in CogBench.
Time Location Character Character Relationship Event Event Relationship Next Moment Event Mental State
CoR 47 179 106 263 701 425 107 417
QA 86 220 162 317 658 402 135 597
CogBench consists of 251 semantically-rich images with a total of 2670 entities, 2245 CoRs, 251 descriptions and 2577 questions, indicating the content contained in each image is complex, showcased in Table 1. The number of CoRs in event-related reasoning and [Mental State Reasoning] is large, which is a manifestation of the rich interesting stories in the images.
3 Experiments
In this section, we will evaluate selected LVLMs, detail evaluation metrics and discuss the results.
3.1 Large Vision-Language Models
We evaluated a selection of recent representative open-source and closed-source LVLMs in this paper, including InstructBLIP-7B [20], Qwen-VL-Chat [21], LLaVA-v1.5-7B and LLaVA-v1.5-13B [22], mPLUG-Owl-2 [6], ShareGPT4V-7B and ShareGPT4V-13B [23], GPT-4V and GPT-4o [1]. A brief introduction of these models is in Appendix D. These open-source models are deployed on a machine with four GeForce RTX 2080 Ti GPUs and a machine with one GeForce RTX 3090 GPU.
3.2 CogBench Evaluation Strategy
Evaluation for Description Task
Here we consider model performance at two levels: low-level Recognition ability and high-level Cognition ability. Evaluation metrics for both levels are calculated using recall scores, referred to as Recognition Score and Cognition Score, respectively.
The Recognition Score is calculated as the ratio of the number of recognized [Entities] to the number of annotated [Entities] in all images. First, we use SpaCy [^6] https://spacy.io/ to extract nouns from the model generated description, and then calculate cosine similarity between embeddings [^7] Implemented with sentence-transformers package (https://www.sbert.net) and all-mpnet-base-v2 is adopted as the model to encode [Entities] and nouns. of annotated [Entities] and extracted nouns. For each entity, if the cosine similarity score between the entity and any noun is greater than a threshold (0.6 in this paper), we consider the entity is recognized by the model.
For Cognition Score, we calculate both scores for each of the eight cognitive reasoning abilities and an overall score using GPT-4 [1]. To enhance objectivity and granularity, GPT-4 is utilized for a binary classification task to assess if generated descriptions include the semantics of annotated CoRs, as detailed in Appendix C. The CoR scores for each type of reasoning capability are then used to compute a recall score for each respective type. The overall Cognition Score is the sum of all CoR scores divided by the total number of CoRs.
Evaluation for VQA Task
For Multiple-Choice Questions in VQA task, we use accuracy as the evaluation metric. As questions are generated based on CoRs, we can both calculate the accuracy for each reasoning capability and the overall cognition capability.
3.3 Results of Description Task
We prompt the selected LVLMs with the following instruction to obtain detailed descriptions about images in CogBench: ‘‘Describe this image in detail.’’ We evaluate the performance of LVLMs on Description task in terms of both recognition and cognition ability. As a reference, we also calculated traditional image captioning evaluation metrics by comparing annotated reference description and model-generated description, and details are shown in Appendix E.
3.3.1 Recognition
Table 2: Recognition score of LVLMs on Description task. For reference, the Recognition Score of Oracle is calculated based on the annotated [Description] in CogBench as an estimated upper bound.
Model Recognition Score
InstructBLIP-7B 0.40
Qwen-VL-Chat 0.42
LLaVA-v1.5-7B 0.40
LLaVA-v1.5-13B 0.41
mPLUG-Owl-2 0.37
ShareGPT4V-7B 0.47
ShareGPT4V-13B 0.49
GPT-4V 0.67
GPT-4o 0.72
Oracle 0.93
Table 2 shows the Recognition Score of models on Description task. GPT-4o achieves the best performance in terms of recognition, which means it can recognize and describe more entities. Besides, GPT-4o and GPT-4V are significantly better than other open-source LVLMs, indicating these open-source LVLMs still have some room for development before reaching the recognition capability of GPT-4. ShareGPT4V-7B and ShareGPT4V-13B achieves better performance than other open-source LVLMs. As it follows the design of LLaVA v1.5, one possible reason is that ShareGPT4V uses a high-quality image-text dataset featuring highly descriptive captions for training, which makes it describe more entities shown in images. GPT-4o, while top-performing, still misses many entities, suggesting room for improvement even on recognition capability.
3.3.2 Cognition
Table 3 shows Cognition Scores of LVLMs on Description task. Similarly, GPT-4o achieves the best performance and there is a large performance gap between GPT-4 and other open-source models. For open-source models, Qwen-VL-Chat achieves the best performance. In terms of different reasoning capabilities, all of the LVLMs show better performance on [Location Reasoning] than others, which is probably because it is a kind of relatively lower-level reasoning. Differently, for [Event Reasoning], [Event Relationship Reasoning], and [Next Moment Event Reasoning], all of the open-source LVLMs show very low performance, indicating they almost do not understand the story in the images at all. The significantly large performance gap on these three kinds of reasoning types between GPT-4 and open-source LVLMs could be a manifestation of the capabilities emerging in GPT-4. Though GPT-4o achieves the best performance, there is also a huge gap between its ability and the Oracle, and the gap is obviously larger than that of recognition scores. This indicates that LVLMs still have a lot of room for development in terms of cognitive abilities.
Table 3: Cognition Scores of LVLMs on Description task evaluated by GPT-4. For reference, the Cognition Score of Oracle is calculated based on the annotated [Description] in CogBench as an estimated upper bound.
Model Time Location Character Character Relationship Event Event Relationship Next Moment Event Mental State Overall
InstructBLIP-7B 0.15 0.54 0.25 0.30 0.10 0.05 0.02 0.17 0.17
Qwen-VL-Chat 0.28 0.57 0.28 0.26 0.14 0.12 0.04 0.13 0.19
LLaVA-v1.5-7B 0.09 0.45 0.15 0.18 0.09 0.05 0.02 0.13 0.13
LLaVA-v1.5-13B 0.13 0.49 0.18 0.24 0.10 0.05 0.04 0.16 0.15
mPLUG-Owl-2 0.06 0.48 0.24 0.21 0.08 0.04 0.02 0.14 0.14
ShareGPT4V-7B 0.19 0.60 0.21 0.22 0.10 0.04 0.03 0.15 0.16
ShareGPT4V-13B 0.23 0.57 0.24 0.26 0.12 0.07 0.03 0.14 0.17
GPT-4V 0.43 0.71 0.47 0.24 0.35 0.31 0.10 0.46 0.37
GPT-4o 0.51 0.77 0.55 0.37 0.44 0.39 0.15 0.40 0.43
Oracle 0.92 0.98 0.94 0.81 0.98 0.92 0.90 0.92 0.93
3.3.3 Case Study of Description Task
Figure 5 shows a case of Description task. In terms of recognition, GPT-4V shows a good performance by recognizing most annotated entities such as boy, girl, man, woman, table, dress, hat, suit and only fails to recognize sofa, television, handbag . LLaVA-v1.5-13B obviously recognizes fewer entities than GPT-4V, as it only recognizes boy, woman, dress. However, GPT-4V fails to understand the story in the image and gets a 0 in terms of cognition. This is because it neither clarifies the location nor the relationships between the characters, nor does it accurately depict the events occurring in the image, etc. LLaVA-v1.5-13B is slightly better and successfully inferred that the woman is the mother. This case demonstrates that CogBench reveals current LVLMs falling short in recognition and cognition, with a gap remaining between their cognitive abilities and human levels.
Figure 5: Case study of the Description task. A representative open-source LVLM LLaVA-v1.5-13B (red frame) and GPT-4V (green frame) are selected for analysis. Recognized entities are marked in blue, and CoRs are marked in green.
3.3.4 Effectiveness Analysis of GPT-based Cognition Evaluation of Description Task
Table 4: CoR accuracy of evaluation methods.
Model Accuracy
ROUGE 0.656
BERTScore 0.635
BLEURT 0.620
DeBERTa 0.693
DocNLI 0.714
GPT-3.5 0.807
GPT-4 0.833
To validate GPT-based cognition evaluation, we manually scored the CoRs of 20 images with a binary scale (0/1) and compared various evaluation methods’ accuracy on this subset. Table 4 reveals that GPT-4 offers the highest accuracy, demonstrating that GPT-based evaluation aligns well with human assessment. Therefore it is effective to assess LVLMs’ performance on the Description task. Implementation details of evaluation methods beyond ChatGPT and GPT-4 can be found in Appendix F.
3.4 Results of VQA Task
Table 5: Model performance on VQA task. Each QA contains four options, with a chance rate of 25%.
Model Time Location Character Character Relationship Event Event Relationship Next Moment Event Mental State Overall
InstructBLIP-7B 0.56 0.65 0.49 0.48 0.36 0.38 0.44 0.43 0.44
Qwen-VL-Chat 0.65 0.82 0.60 0.54 0.51 0.45 0.47 0.51 0.54
LLaVA-V1.5-7B 0.58 0.81 0.54 0.55 0.46 0.45 0.54 0.53 0.53
LLaVA-V1.5-13B 0.70 0.82 0.65 0.60 0.50 0.47 0.58 0.57 0.57
mPLUG-Owl-2 0.51 0.82 0.59 0.55 0.46 0.48 0.47 0.52 0.53
ShareGPT4V-7B 0.58 0.80 0.64 0.54 0.49 0.40 0.51 0.54 0.53
ShareGPT4V-13B 0.67 0.80 0.65 0.56 0.49 0.50 0.6 0.55 0.56
GPT-4V 0.66 0.84 0.73 0.62 0.63 0.63 0.70 0.69 0.67
GPT-4o 0.78 0.91 0.77 0.73 0.75 0.73 0.83 0.77 0.77
Human 0.99 0.96 0.99 0.94 0.96 0.96 0.96 0.93 0.95
Table 5 shows the performance of LVLMs on VQA task. Consistent with results in previous sections, GPT-4o achieves the best performance and there is a performance gap between GPT-4o and GPT-4V. There is also a performance gap between GPT-4 and open-source models. For open-source models, ShareGPT4V-13B and LLaVA-v1.5-13B achieves better performance than other LVLMs based on 7B LLMs, which shows the importance of LLM size to LVLMs on this task. As for LVLMs based on 7B LLMs, the performance of InstructBLIP is the worst and there is also a gap between InstructBLIP and other models. One possible reason is that it only finetunes the Q-former for instruction-tuning as Q-former has limited capacity compared with LLMs. The performance of other models are similar. Consistent with previous findings, reasoning about location is also the easiest for LVLMs and event-related reasoning are more difficult. There is also a large gap between the performance of LVLMs and human beings. Note that the accuracy of Human in Table 5 is calculated based on the responses of five healthy people. They all have obtained a bachelor’s degree and are between the ages of 20 and 30.
4 Related Work
Evaluation Benchmark for LVLMs.
To better understand emerging capabilities of LVLMs, many scholars have proposed different evaluation benchmarks. LVLM-eHub [24] evaluates LVLMs’ multimodal capabilities across six categories using various publicly available computer vision datasets. MME [10], MMBench [11] and SEED Bench [12] use True/False Questions or Multiple-Choice Questions to evaluate different abilities of LVLMs. MM-VET [25] evaluates LVLMs in terms of their integrated VL capabilities via open-ended questions. Different from them, CogBench mainly focus on high-level cognition evaluation of LVLMs. Though some of them also consider cognition as one of the evaluation dimensions, most images they use can only be used to evaluate limited aspects of reasoning.
Image Captioning.
Image Captioning is a classical VL task. As model capabilities advance, researchers strive to enhance their ability to describe images in detail. Krause et al. [26] propose Image Paragraph Captioning, tasking models with generating a descriptive paragraph for an image. Recently, some researchers [17, 18, 19, 23] are trying to leveraging the ability of LLMs to generate more detailed image descriptions. None of these tasks consider to evaluate the high-level cognitive ability of models through description. HL dataset [27] requires models to generate high-level captions, but it only considers three aspects (scene, action, rationale). The content of most images in existing datasets does not reach the level of a story. This reveals the need for higher-level datasets like CogBench.
Visual Reasoning.
Visual Reasoning is closely related to the cognitive ability of models. VCR [28] tasks models with answering visual questions using commonsense reasoning and justifying their answers. VisualCOMET [29] is a framework of visual commonsense reasoning tasks to predict past, future events, and present intents. Hessel et al. [30] utilize images from VCR and Visual Genome [31] to evaluate the ability of models to perform abductive reasoning. Fu et al. [32] propose a task to identify the time and location of a given image. CURE [33] is proposed to measure both the zero-shot reasoning performance and consistency of VLMs. Similarly, compared CogBench, these tasks considers less kinds of reasoning and CogBench can be seen as the next step of these efforts.
5 Conclusion
In this paper, we incorporated the idea of Cookie Theft picture description task into the evaluation of the high-level cognitive abilities of LVLMs and designed a novel evaluation benchmark called CogBench. Images in CogBench are of high quality and contain rich cognitive reasoning, which makes it different from existing image datasets. Experiments show that there is still a large gap between the cognitive ability of LVLMs and human beings, indicating CogBench is a challenging benchmark.
Appendix A Image Annotation Instruction
Figure 6 shows the image annotation instruction for annotators of CogBench.
You are going to see some pictures. Each picture tells a story and requires different kinds of reasoning to fully understand the story behind it. You will be first asked to identify the entities and reasoning processes in the picture. Then, you will need to describe the story of the picture based on your identified entities and reasoning processes. First, you will be asked to identify the entities in the picture. The annotation format is [A, B, C], where A, B, C are entities. [Entities]: Please list all entities appearing in the picture, including people, animals, objects etc. You are encouraged to list as many entities as possible. Note that these entities need to be in your picture description afterwards. For entities that are difficult to recognize, please do not list them here or describe them. Then, you will be asked to identify different reasoning processes in the picture. The annotation format should follows the structure [A1 + A2 -> B], where A1 and A2 are premises and B is the conclusion. Note that if you write a conclusion, there must be at least one premise. Do not write a conclusion only, like [B]. Please write one conclusion at a time, do not write a reasoning process like [A1->B->C], which should be split into two. Each picture does not necessarily requires all the reasoning. Please write None, if a picture does not involve a specific kind of reasoning or it is not important in the picture. [Special Time Reasoning] Please write your reasoning processes about the special time of the story in the picture, e.g. festivals, seasons etc. The special time is usually relevant to the story of the picture. For instance, if it is daytime in a picture, it is easily recognized and requires no reasoning and there is nothing special, you can write None. However, if there is a lamp on or a clock indicating a specific time, you can write down your reasoning about time. [Location Reasoning] Please write your reasoning processes about the location of the story in the picture, e.g. near the school. [Character Reasoning]: Please write your reasoning processes about the character of the subjects in the picture, e.g. teacher, doctor etc. [Character Relationship Reasoning]: Please write your reasoning processes about relationship between characters in the picture, e.g. friendship. [Event Reasoning]: Please write your reasoning processes about events happened in the current moment and previous moments in the picture based on clues in the picture. Note that you only need to annotate those high-level events and you can ignore those low-level events. For instance, “the woman is looking at the man” is a low level event and you can ignore its reasoning process. Differently, the reasoning process [A mother is busy cooking. + A boy is fetching cookies behind the mom. + A girl is shushing the boy. -> The boy is stealing cookies.] is a reasoning about high-level event stealing and you should write it down. [Event Relationship Reasoning]: Please write your reasoning processes about relationship between different events in the picture. These events are usually linked through causal and temporal relations. Note that events in this part do not necessarily appears in the [Event Reasoning] part as some events here are low-level events. [Next Moment Event Reasoning]: Please write your reasoning processes about the event that will happen in the next moment. Note that you only need to write down events that have a very high probability of happening, instead of guessing what might happen next. [Mental State Reasoning]: Please write your reasoning processes about mental state of subjects in the picture, e.g. daydreaming, happy etc. You need to reason as best you can about the mental states of all the subjects in the picture, unless they are not showing obvious emotions. Finally, you will be asked to describe the picture in as much detail as you can. [Description]: Please describe all you see in the picture in a paragraph based on the entities and reasoning processes you identified above, ensuring that all of them are included in your description. Each picture has a story behind it and you need to tell that story through your description.
Figure 6: Image annotation instruction for annotators.
Appendix B Prompt of CoR-based GPT-assisted Question Generation
Figure 7 shows an example prompt of CoR-based GPT-assisted Question Generation for GPT-4. This prompt is used to generate questions based on [Event Reasoning] CoRs. Prompts for other CoR types are similar to this one.
We have a description of an image and the description tells a detailed story unfolding in the image. In the process of describing an image, it is often necessary to engage in reasoning about events based on the clues within the image, leading to certain conclusions. For example, when we see the wind is blowing outside, and a man is reading a newspaper in the telephone booth, we can infer that he is actually hiding from the wind in the telephone booth. Therefore, in this task, in addition to the image description, the reasoning processes about event within the image description have also been extracted. For each reasoning process, we use A1+A2+…->B to represent it, where A1, A2, … are clues we observed in the picture and B represents the conclusion about event we inferred. Thus, given an image description and the reasoning processes about event, our task is: 1) Generate a question based on reasoning processes about event. 2) Generate four options: A, B, C, and D. There is only one correct answer among the four options, which is consistent with the description and reasoning processes provided. The correct answer option should be randomly chosen from A, B, C, and D. For those incorrect options (distractors), you are encouraged to hallucinate some clues that are highly relevant to the question and the description but do not actually consistent with the description. That is, you can distort the facts in the description and reasoning processes using elements related to the question to generate some easily selectable distractors. It would be better if you can generate some distractors that are similar to but different from the correct option. Please avoid situations where the correct option is significantly longer or shorter than the distractors. For example, if the description is "There are some snow on the ground and it is windy, … We can see it is cold. Inside a phone booth, a man is smiling while looking at a newspaper. He is sheltering from the cold wind in the phone booth…" and the question is "Why can we tell that the man is seeking shelter for warmth?", you can use "newsstand", which is related to "seeking shelter for warmth" in the question, to distort the fact in description "in a phone booth." Then you can get "the man is in the newsstand." Similarly, you can hallucinate a question related distractor "it is raining and a man is smiling and reading a newspaper in a phone booth," which is similar to the correct option "it is windy and a man is smiling and reading a newspaper in a phone booth," but different from it and inconsistent with the description. 3) Generate the the letter corresponding to the correct answer, that is A, B, C, or D. Here are some examples: ———- [Description]: There are some snow on the ground and it is windy, indicating it is winter. There are two men and two women standing on the roadside. There is a sign that says "NO STANDING BUS STOP", indicating it is near a bus stop. A man is standing on the road side, wrapping his coat tightly around himself, and peering out onto the road. They are probably waiting for a bus here. We can see it is cold. Inside a phone booth, a man is smiling while looking at a newspaper. He is sheltering from the cold wind in the phone booth. He looks happy, because it is warm there. Two women are also wrapping their coats tightly and looking at the man in the phone booth. They are probably friends and standing together. They are unhappy with the man. There are some buildings by the road. [Event Reasoning]: It is windy and cold. + A man is standing in a phone booth reading newspaper. -> The man is sheltering from the cold wind in the phone booth. [Generated Multiple-Choice Questions]: What is the man doing in the phone booth? A. Making a phone call. B. Reading a book. C. Avoiding someone he doesn’t want to see. D. Sheltering from the cold wind. Correct Answer: [D] Why can we tell that the man is seeking shelter for warmth? A. It is windy and a man is smiling and reading a newspaper in a newsstand. B. It is raining and a man is smiling and reading a newspaper in a newsstand. C. It is windy and a man is smiling and reading a newspaper in a phone booth. D. It is raining and a man is smiling and reading a newspaper in a phone booth. Correct Answer: [C] ———- Please: 1). Generate at least one question for each reasoning process. 2). Generate more diverse questions, try to generat questions from different perspectives or angles and don’t limit yourself to the question templates provided in the examples. 3). Avoid generating repetitive questions with similar meanings.
Figure 7: An example prompt of CoR-based GPT-assisted Question Generation for GPT-4 to generate questions based on [Event Reasoning].
Appendix C Prompt of GPT-based Description Cognition Evaluation
Figure 8 and Figure 9 shows prompts of ChatGPT or GPT-4 for cognition evaluation of Description task.
For CoR types other than [Event Relationship Reasoning], we task GPT-4 with determining whether the conclusion in each CoR is mentioned in the description. The prompt is shown in Figure 8. For [Event Relationship Reasoning], we task GPT-4 with determining whether the causal relationship between events (i.e. the whole CoR), as annotated, is present in the description. The prompt is shown in Figure 9.
Given a <DESCRIPTION> and some <KEY POINT>s, please tell me if the <DESCRIPTION> explicitly presents the exact or similar semantics of each <KEY POINT>. The following points are required:
1) Instead of reasoning about whether each <KEY POINT> is possibly correct based on the <DESCRIPTION>, you only need to determine whether the <DESCRIPTION> mentions the semantics in the <KEY POINT>. 2) Do not overlook the semantics in the <DESCRIPTION> that are semantically equivalent to the <KEY POINT> but expressed in different ways. For instance, if the <DESCRIPTION> mentions "The woman is playing with her son…", we can tell it successfully includes semantics in the <KEY POINT> "The woman is the mother of the boy." 3) If several possible scenarios are listed using ’or’ at a <KEY POINT>, you only need to determine whether one of these scenarios is mentioned in the <DESCRIPTION>. Assign a score of 0 or 1 to each <KEY POINT>, where 0 represents NO and 1 represents YES. <DESCRIPTION>: {Description generated by models.} <KEY POINT>: 1. {Annotated key point 1.} 2. {Annotated key point 2.} N. {Annotated key point N.} Please write your answers in “[ ]” with 0 or 1 in the following format (number + square brackets): 1. [1] 2. [0] Your answers to the {N} <KEY POINT>(s) above: 1. [ ] 2. [ ] … N. [ ]
Figure 8: Evaluation prompt for GPT-4 of Reasoning types other than [Event Relationship Reasoning].
Given a <DESCRIPTION> and some <EVENT RELATIONSHIP>s, please tell me whether this <DESCRIPTION> clearly depicts the cause-and-effect relationships between events. The format of a <EVENT RELATIONSHIP> follows the structure "A1 + A2 + … + An -> B", where A1, A2, …, An and B are events. Events A1, A2, …, An are the causes of event B, and event B is the result caused by events A1, A2, …, An. The criteria for judgment lie in whether the <DESCRIPTION> mentions these events and clearly depicts the causal relationships between them. Assign a score of 0 or 1 to each <EVENT RELATIONSHIP>, where 0 represents NO and 1 represents YES. <DESCRIPTION>: {Description generated by models.} <EVENT RELATIONSHIP>: 1. {Annotated event relationship 1.} 2. {Annotated event relationship 2.} N. {Annotated event relationship N.} Please write your answers in “[ ]” with 0 or 1 in the following format (number + square brackets): 1. [1] 2. [0] Your answers to the {N} <EVENT RELATIONSHIP>(s) above: 1. [ ] 2. [ ] … N. [ ]
Figure 9: Evaluation prompt for GPT-4 of [Event Relationship Reasoning].
Appendix D Introduction to Selected LVLMs
Table 6: LVLMs evaluated in this paper.
Model Visual Encoder Language Model
InstructBLIP-7B EVA-CLIP ViT-g Vicuna-7B
Qwen-VL-Chat OpenCLIP ViT-G Qwen-7B
LLaVA-v1.5-7B CLIP ViT-L Vicuna-7B
LLaVA-v1.5-13B CLIP ViT-L Vicuna-13B
mPLUG-Owl-2 CLIP ViT-L LLaMA-7B
ShareGPT4V-7B CLIP ViT-L Vicuna-7B
ShareGPT4V-13B CLIP ViT-L Vicuna-13B
GPT-4V
GPT-4o
  • InstructBLIP [20] builds upon BLIP-2 [34]. It consists of an image encoder, a LLM, and a Query Transformer (Q-Former). During instruction tuning, only the Q-Former is updated. We use “blip2-vicuna-instruct” for testing.
  • Qwen-VL-Chat [21] is a instruction-tuned VL chatbot based on Qwen-VL. Its training process consists of two stages of pre-training and a final stage of instruction fine-tuning. As for architecture, it consists of a vision encoder, a LLM, and position-aware vision-language adapter. We test “Qwen-VL-Chat”.
  • LLaVA v1.5 [22] is an upgraded version of LLaVA [5]. LLaVA connects a vision encoder and LLM for general-purpose visual and language understanding. It is instruction-tuned on the language-image instruction-following data generated by language-only GPT-4 [1]. By using CLIP-ViT-L-336px with an MLP projection and adding academic-task-oriented VQA data with simple response formatting prompts, LLaVA v1.5 achieves better performance. “llava-v1.5-7b” and “llava-v1.5-13b” are tested.
  • mPLUG-Owl-2 [6] mainly comprises a fundamental vision encoder, a visual abstractor, and a language decoder. It also adopts a two-stage training strategy, comprising pre-training and visual instruction tuning. We test “mplug-owl2-llama2-7b”.
  • ShareGPT4V [23] follows the design of LLaVA v1.5. They incorporate a large-scale resource featuring highly descriptive captions into both the pre-training and supervised fine-tuning phases of ShareGPT4V model. We test“ShareGPT4V-7B” and “ShareGPT4V-13B”.
  • GPT-4V [1] is one of the most powerful LVLMs in the world developed by OpenAI. The version of “gpt-4-vision-preview” is tested.
  • GPT-4o [1] is currently OpenAI’s latest and most powerful multimodal model. It is a single model trained end-to-end across text, vision, and audio. The version of “gpt-4o” is tested.
Table 6 shows an overview of the design of different LVLMs.
Appendix E Evaluation Results of Description Task based on Traditional Image Captioning Metrics
Table 7: Model performance on Description task evaluated by traditional image captioning evaluation metrics.
Model METEOR CIDEr BLEU-1 BLEU-2 BLEU-3 BLEU-4
InstructBLIP-7B 0.130 0.043 0.255 0.127 0.063 0.033
Qwen-VL-Chat 0.128 0.036 0.232 0.121 0.059 0.031
LLaVA-V1.5-7B 0.146 0.054 0.309 0.158 0.076 0.037
LLaVA-V1.5-13B 0.146 0.051 0.312 0.160 0.076 0.037
mPLUG-Owl-2 0.132 0.035 0.260 0.126 0.057 0.027
ShareGPT4V-7B 0.162 0.017 0.259 0.120 0.050 0.024
ShareGPT4V-13B 0.165 0.024 0.278 0.129 0.055 0.026
GPT-4V 0.208 0.001 0.234 0.115 0.052 0.025
GPT-4o 0.210 0.008 0.254 0.128 0.061 0.030
Table 7 shows the performance of models on traditional image captioning evaluation metrics. Following Krause et al. [26], we use METEOR [35], CIDEr [36], BLEU-1, BLEU-2, BLEU-3, and BLEU-4 [37] to evaluate the performance of models on Description task. Similar to the findings of Zhu et al. [18], it can be observed that traditional image captioning evaluation metrics are not quite suitable for evaluating the Description task. There are two possible reasons. The first possible reason is that image descriptions are longer and more flexible than traditional image captions. The second possible reason is that Description task requires evalution metrics to consider high-level semantics in description, while traditional image captioning evaluation metrics only concerns low-level information.
Appendix F Implementation of Non-GPT-based Cognition Evaluation Methods of Description Task
The task of cognition evaluation of Description task is to determine whether a CoR is mentioned in the description. Apart from ChatGPT or GPT-4, some other evaluation methods are implemented to perform this classification task, as shown in Table 4.
For methods based on ROUGE [38], BERTScore [39], and BLEURT [40], we first split the description into sentences, then use a CoR as a reference to calculate the (recall) score for each sentence compared to the CoR. Then, the highest score among all calculated scores is taken as the score of the CoR corresponding to the description. Finally, the score is converted into 0/1 using a threshold.
We also tried Natural Language Inference (NLI) models to perform the task. First, we use DeBERTa [41] to perform sentence-level NLI task similar to methods mentioned above. If there is at least one “Entailment” for all the sentences, the score of the CoR will be 1. The model we adopted is mDeBERTa-v3-base-xnli-multilingual-nli-2mil7. The second NLI model we tried is DocNLI [42], which can directly take the description and CoR as input and do the classification task.
Appendix G Ethical Considerations
Most images in CogBench are manually collected from Pinterest. We follow the terms of service of Pinterest to collect these images. The images are used as fair use for research purposes only. We will share the data with other researchers who will follow the ethical considerations as established in this study. We have compensated our annotators at a rate of \$0.60 per image, which aligns with the university’s recruitment standards for labeling personnel. This amount is offered as an appropriate token of gratitude for their valuable time and efforts in assisting with our project. We answer any questions our annotators have during their annotation process promptly. They are free to take breaks or quit the annotation task at any time.